Sulfur hexafluoride, SF6 (g), serves as a chemical insulator. A 5.0 L sample of this gas was collected at 205.0 ° C and 350 kPa. What pressure must be applied to the gas sample to reduce its volume to 1.7 L if the temperature is 25 ° C?


1
Expert's answer
2021-05-18T07:52:41-0400

"V_1 = 5.0 \;L \\

T_1 = 205.0 + 273 = 478 \;K \\

P_1 = 350 \;kPa \\

V_2 = 1.7 \; L \\

T_2 = 25 + 273 = 298 \;K \\

\frac{P_1 \times V_1}{T_1}= \frac{P_2 \times V_2}{T_2} \\

P_2 = \frac{P_1 \times V_1 \times T_2}{T_1 \times V_2} \\

= \frac{350 \times 5.0 \times 298}{478 \times 1.7} \\

= 641.7 \;kPa"

Answer: 641.7 kPa